# TGIF: Text-Guided Inpainting Forgery Dataset

### v1

@@ IV-A Experimental Setup
We performed the evaluation on the test set, separating the spliced subsets (SD2-sp and PS-sp) and fully regenerated subsets (SD2-fr and SDXL-fr). Additionally, to make the detection more challenging, we evaluated the performance on compressed variants of the real and fake images. That is, we used the JPEG and WEBP compression standards, using quality factors of 80 and 60.
For image forgery localization we use the mean pixel-level F1 score of the fake images as a performance metric, which is the harmonic mean of the precision and recall. We utilize a threshold of 0.5 to decide whether pixels are real or fake. This threshold demonstrates the calibration of the models, which is essential when deploying the models in the wild and requiring human interpretation. We selected several recent high-performing AI-based IFL methods, namely the PSCC-Net [19], SPAN [20], ImageForensicsOSN [21], MVSS-Net++ [12], Mantranet [22], CAT-Net (v2) [23], TruFor [8], and MMFusion [13].
For synthetic image detection, as a performance metric, we use the AUC score. We opted to report the threshold-agnostic AUC rather than threshold-specific image-level F1 score, because several SID methods reported high F1 scores while actually demonstrating bad performance. More specifically, these methods often classified both real and fake images as fake, which, due to the class imbalance, significantly boosted the F1 score. Since the AUC represents the trade-off between the false positive rate and the false negative rate for a range of thresholds, it better reflects the detection performance, in this case. We selected several AI-based SID methods available in the SIDBench framework [47], namely CNNDetect [24], DIMD [31], Dire [32], FreqDetect [25], UnivFD [26], Fusing [48], GramNet [27], LGrad [28], NPR [29], RINE [33], DeFake [34], and PatchCraft [30]. Some methods provide multiple models, trained on different data. We only show the best performing in this paper. Although some of these methods were trained on GANs, some of them have demonstrated generalization performance to DMs. Hence, also methods trained on GANs were included in our evaluation set.
@@ IV-B Image Forgery Localization Methods
Table I reports the F1 values for all evaluated IFL methods. We highlight F1 values above 0.7. This threshold was chosen arbitrary, but allows to quickly grasp which methods demonstrate good detection performance. For the splicing datasets SD2 and PS, we observe that only CAT-Net, TruFor, and MMFusion report high F1 scores. Additionally, Mantranet only reports a mediocre F1 value for the PS dataset, while it has low performance for SD2. All other methods demonstrate low detection performance.
For the fully regenerated datasets SD2-fr and SDXL-fr, no method is able to detect and localize these manipulations. This is as expected, since regenerating the image removes most invisible traces such as compression artifacts and camera noise. In fact, diffusing and regenerating an image is an existing attack against forgery detection methods [49]. This highlights the threat of local image manipulations using diffusion models that fully regenerate images; these cannot be detected by traditional image forgery localization methods. This threat is enlarged when one considers the widespread use of GenAI models in applications such as super-resolution[50] and image compression [51]. Although these applications can be used with harmless intentions, they may remove traces used by traditional image forgery localization methods.
We only analyzed the influence of compression for the best performing methods in the uncompressed case, and only for the splicing datasets. When performing JPEG compression, F1 scores drop significantly. Moreover, this performance drop is amplified when performing WEBP compression. Hence, this demonstrates a need to develop IFL methods that are robust to compression.
TABLE I: Evaluation of Image Forgery Localization Methods on our TGIF Dataset (F1 Score).
@@ IV-C Synthetic Image Detection Methods
Method Trained on Uncompressed JPEG Q80 JPEG Q60 WEBP Q80 WEBP Q80
SD2-fr SDXL-fr SD2-sp PS-sp SD2-fr SDXL-fr SD2-fr SDXL-fr SD2-fr SDXL-fr SD2-fr SDXL-fr
CNNDetect [24] ProGAN 0.57 0.61 0.52 0.51 0.57 0.62 0.56 0.61 0.55 0.60 0.55 0.60
DIMD [31] LDM 1.00 0.94 0.66 0.54 1.00 0.95 0.99 0.92 0.96 0.71 0.90 0.61
Dire [32] ADM 0.49 0.62 0.49 0.51 0.49 0.61 0.49 0.60 0.50 0.60 0.48 0.60
FreqDetect [25] GANs 0.50 0.40 0.53 0.56 0.72 0.71 0.71 0.72 0.66 0.61 0.62 0.62
UnivFD [26] ProGAN 0.82 0.8 0.62 0.63 0.73 0.78 0.70 0.73 0.71 0.74 0.69 0.73
Fusing [48] ProGAN 0.55 0.47 0.52 0.54 0.52 0.59 0.53 0.61 0.50 0.53 0.50 0.54
GramNet [27] GANs 0.82 0.55 0.57 0.63 0.48 0.56 0.47 0.54 0.49 0.54 0.50 0.53
LGrad [28] ProGAN 0.85 0.83 0.55 0.60 0.51 0.57 0.51 0.64 0.47 0.42 0.47 0.41
NPR [29] ProGAN 0.51 0.67 0.50 0.51 0.50 0.30 0.51 0.68 0.49 0.69 0.49 0.31
RINE [33] ProGAN 0.89 0.89 0.64 0.65 0.79 0.88 0.75 0.85 0.77 0.85 0.76 0.87
LDM 0.97 0.93 0.61 0.64 0.84 0.89 0.80 0.82 0.75 0.76 0.72 0.77
DeFake [34] Diffusion 0.65 0.61 0.55 0.51 0.54 0.45 0.54 0.43 0.54 0.45 0.53 0.49
PatchCraft [30] ProGAN 0.98 0.95 0.55 0.57 0.66 0.67 0.60 0.58 0.60 0.57 0.59 0.61
Table II shows the performance (AUC) results for all evaluated SID methods. We highlight AUC values above 0.8. This threshold was chosen arbitrarily, but allows to quickly grasp which methods demonstrate good detection performance. For the spliced datasets SD2-sp and PS-sp, no SID method is able to detect the synthetic spliced regions. For the fully regenerated datasets, we observe that 6 out of 12 methods are able to detect synthetic images in uncompressed form. Surprisingly, some methods trained on GANs (such as UnivFD, LGrad and RINE) demonstrate better performance than methods trained on DMs (such as DeFake). Additionally, only two of those methods survive JPEG compression, and only one survives WEBP compression. Moreover, some methods (such as DIMD) are more robust in detecting SD2 than SDXL, whereas others (such as RINE) are more robust in detecting SDXL than SD2. No single method is robust for both SD2 and SDXL on all evaluated compression levels. This demonstrates that these methods are complementary.
In summary, we demonstrated that several state-of-the-art SID methods are able to detect fully regenerated inpainted images by SD2 and SDXL. However, they fail to detect the spliced inpainted images (even though the spliced regions were generated synthetically). Additionally, their performance decreases when exposed to compression, with a notable difference between WEBP and JPEG compression. Finally, the tested SID methods are not able to localize the inpainted area, but rather produce a SID score for the whole image.
@@ V Discussion & Conclusion
Our benchmark analysis shows that some of the existing IFL methods are able to detect and localize spliced images, whereas they fail to localize the inpainted area in fully regenerated images. In contrast, some of the existing SID methods are able to detect fully regenerated images, yet lack the ability to localize the synthetic inpainted area. Additionally, they cannot detect the synthetic data when it is spliced into the image. These limitations highlight the need for new forensic methods, perhaps combining aspects from both IFL and SID methods.
Furthermore, we showed that forensic methods suffer from a notable decrease in performance when exposed to compression, particularly with WEBP. This highlights the need to include compressed images during training.

### v2

@@ IV-A Experimental Setup
We performed the evaluation on the test set, separating the spliced subsets (SD2-sp and PS-sp) and fully regenerated subsets (SD2-fr and SDXL-fr). Additionally, to make the task more challenging, we evaluated the performance on compressed variants of the real and fake images. That is, we used the JPEG and WEBP compression standards, using quality factors of 80 and 60.
For IFL, we use the mean pixel-level F1 score of the fake images as the performance metric, which is the harmonic mean of the precision and recall. We utilize a threshold of 0.5 to decide whether pixels are real or fake. This threshold assesses the calibration of the models, which is essential when deploying the models in the wild and requiring human interpretation. We selected several recent high-performing AI-based IFL methods, namely the PSCC-Net [19], SPAN [20], ImageForensicsOSN [21], MVSS-Net++ [12], Mantranet [22], CAT-Net (v2) [23], TruFor [8], and MMFusion [13].
For SID, we use the AUC score as the performance metric. We opted to report the threshold-agnostic AUC rather than threshold-specific image-level F1 score, because several SID methods reported high F1 scores while actually demonstrating bad performance. More specifically, these methods often classified both real and fake images as fake, which, due to the class imbalance, significantly boosted the F1 score. Since the AUC represents the trade-off between the false positive and false negative rate for a range of thresholds, it better reflects detection performance, in this case. We selected several AI-based SID methods available in the SIDBench framework [47], namely CNNDetect [24], DIMD [31], Dire [32], FreqDetect [25], UnivFD [26], Fusing [48], GramNet [27], LGrad [28], NPR [29], RINE [33], DeFake [34], and PatchCraft [30]. Some methods provide multiple models, trained on different data. We only show the best performing in this paper. Although some of these methods were trained on GANs, some of them have demonstrated generalization performance to DMs. Hence, also methods trained on GANs were included in our evaluation set.
@@ IV-B Image Forgery Localization Methods
Table I reports the F1 values for all evaluated IFL methods. We highlight F1 values above 0.7. This threshold was chosen arbitrarily, but allows to quickly grasp which methods demonstrate good detection performance. For the splicing datasets SD2 and PS, we observe that only CAT-Net, TruFor, and MMFusion report high F1 scores. Additionally, Mantranet only reports a mediocre F1 value for the PS dataset, while it has low performance for SD2. All other methods demonstrate low detection performance.
For the fully regenerated datasets SD2-fr and SDXL-fr, no method is able to detect and localize these manipulations. This is expected, since regenerating the image removes most invisible traces such as compression artifacts and camera noise. In fact, diffusing and regenerating an image is an existing attack against forgery detection methods [49]. This highlights the threat of local image manipulations using diffusion models that fully regenerate images; these cannot be detected by traditional image forgery localization methods. This threat is enlarged when one considers the widespread use of generative models in applications such as super-resolution [50] and image compression [51]. Although these applications can be used with harmless intentions, they may remove traces used by traditional IFL methods.
We only analyzed the influence of compression for the best performing methods in the uncompressed case, and only for the splicing datasets. When performing JPEG compression, F1 scores drop significantly. Moreover, this performance drop is amplified when performing WEBP compression. This demonstrates a need to develop IFL methods that are robust to compression.
TABLE I: Evaluation of Image Forgery Localization Methods on our TGIF Dataset (F1 Score). We only evaluated compressed images on IFL methods and sub-datasets exhibiting decent performance on uncompressed images.
@@ IV-C Synthetic Image Detection Methods
Method Trained on Uncompressed JPEG Q80 JPEG Q60 WEBP Q80 WEBP Q60
SD2-fr SDXL-fr SD2-fr SDXL-fr SD2-fr SDXL-fr SD2-fr SDXL-fr SD2-fr SDXL-fr
CNNDetect [24] ProGAN 0.57 0.61 0.57 0.62 0.56 0.61 0.55 0.60 0.55 0.60
DIMD [31] LDM 1.00 0.94 1.00 0.95 0.99 0.92 0.96 0.71 0.90 0.61
Dire [32] ADM 0.49 0.62 0.49 0.61 0.49 0.60 0.50 0.60 0.48 0.60
FreqDetect [25] GANs 0.50 0.40 0.72 0.71 0.71 0.72 0.66 0.61 0.62 0.62
UnivFD [26] ProGAN 0.82 0.8 0.73 0.78 0.70 0.73 0.71 0.74 0.69 0.73
Fusing [48] ProGAN 0.55 0.47 0.52 0.59 0.53 0.61 0.50 0.53 0.50 0.54
GramNet [27] GANs 0.82 0.55 0.48 0.56 0.47 0.54 0.49 0.54 0.50 0.53
LGrad [28] ProGAN 0.85 0.83 0.51 0.57 0.51 0.64 0.47 0.42 0.47 0.41
NPR [29] ProGAN 0.51 0.67 0.50 0.30 0.51 0.68 0.49 0.69 0.49 0.31
RINE [33] ProGAN 0.89 0.89 0.79 0.88 0.75 0.85 0.77 0.85 0.76 0.87
LDM 0.97 0.93 0.84 0.89 0.80 0.82 0.75 0.76 0.72 0.77
DeFake [34] Diffusion 0.65 0.61 0.54 0.45 0.54 0.43 0.54 0.45 0.53 0.49
PatchCraft [30] ProGAN 0.98 0.95 0.66 0.67 0.60 0.58 0.60 0.57 0.59 0.61
Table II shows the performance (AUC) results for all evaluated SID methods. We highlight AUC values above 0.8. This threshold was chosen arbitrarily, but allows to quickly grasp which methods demonstrate decent detection performance. Note that we do not include the performance results for the spliced datasets (SD2-sp and PS-sp), since we cannot expect SID to detect these types of local manipulation, as they were not designed for this purpose. We have verified that, indeed, no SID method is able to detect the synthetic spliced regions.
We observe that 6 out of 12 methods are able to detect (fully regenerated) synthetic images in uncompressed form. Surprisingly, some methods trained on GANs (such as UnivFD, LGrad and RINE) demonstrate better performance than methods trained on DMs (such as DeFake). Additionally, only two of those methods survive JPEG compression, and only one survives WEBP compression. Moreover, some methods (such as DIMD) are more robust in detecting SD2 than SDXL, whereas others (such as RINE) are more robust in detecting SDXL than SD2. No single method is robust for both SD2 and SDXL on all evaluated compression levels. This demonstrates that these methods are complementary.
In summary, we demonstrated that several state-of-the-art SID methods are able to detect fully regenerated inpainted images by SD2 and SDXL. However, they are not able to localize the inpainted area, but rather produce a SID score for the whole image. This is expected as they were not designed to detect the spliced inpainted images (even though the spliced regions were generated synthetically). Finally, their performance decreases when exposed to compression, with a notable difference between WEBP and JPEG compression.
@@ V Discussion & Conclusion
Our benchmark analysis shows that some of the existing IFL methods are able to detect and localize spliced images, whereas they fail to localize the inpainted area in fully regenerated images. In contrast, some of the existing SID methods are able to detect fully regenerated images, yet lack the ability to localize the synthetic inpainted area. These limitations highlight the need for new forensic methods, leveraging elements from both IFL and SID methods. Furthermore, we showed that forensic methods suffer from a notable decrease in performance when exposed to compression, particularly with WEBP. This highlights the need to include compressed images during training.